OpenStudy (anonymous):

evaluate y(x-5) for x=6 and y=4 y(x-5)=

jimthompson5910 (jim_thompson5910):

replace y with 4 and x with 6 to get y*(x-5) 4*(6-5) ???

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

i got a x for it show me 4

jimthompson5910 (jim_thompson5910):

not sure what you mean

OpenStudy (anonymous):

4(6-5) = 4(1) = 4

OpenStudy (anonymous):

i lost it turn red on me

OpenStudy (anonymous):

evaluate v(u-3) for u=8 and v=2 v(u-3)=

jimthompson5910 (jim_thompson5910):

plug in u = 8 and v = 2, then evaluate

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

z-7,z=17

jimthompson5910 (jim_thompson5910):

tell me what you got

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

translate the plrase to an algebraic expression state what each variable represents. five more then a numbler

jimthompson5910 (jim_thompson5910):

you got what?

OpenStudy (anonymous):

im stuck on it

OpenStudy (anonymous):

x+5

OpenStudy (anonymous):

x is the number, +5 more than it

OpenStudy (anonymous):

translate the plrase to an algebraic expression state what each variable represents. one-fifth of a number plus one-fifth of a different number